A person estimates that he can afford a mortgage payment of $1,200 per month. He can obtain a 30-year mortgage at an interest rate of 11.25 percent. What is the largest mortgage loan he can afford?
1
Expert's answer
2018-02-15T08:18:07-0500
If a person estimates that he can afford a mortgage payment of $1,200 per month and can obtain a 30-year mortgage at an interest rate of 11.25 percent, then the total payment is PV = 1,200*12*30 = $432,000. So, the largest mortgage loan he can afford is:
FV*1.1125^30 = 432,000,
FV = $17,639.37.